OpenStudy (rissa):

Find the constant of variation k for the direct variation. 4x = -6y

jimthompson5910 (jim_thompson5910):

4x = -6y -6y = 4x y = (4x)/(-6) y = -(4/6)x y = (-2/3)x Now the equation is in the form y = kx where k = -2/3